OpenStudy (anonymous):

Which of the following tables represents a proportional relationship? x 2 4 6 8 y 3 4 5 6 x 2 3 4 5 y 3 6 9 12 x 3 5 7 9 y 4.5 7.5 10.5 13.5 x 1 4 7 10 y 0 3 6 9

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

The tables must be equal to y = kx.

OpenStudy (igreen):

Take two points from the first table, and see if k is the same in both of them. (2, 3) (4, 4) y = kx (2, 3) 3 = k(2) 3 = 2k 3/2 = k (4, 4) 4 = k(4) 4 = 4k 1 = k This table is NOT proportional because it has different constants(k).

OpenStudy (igreen):

Now take two points from the second table: (2, 3) (3, 6) y = kx (2, 3) 3 = k(2) 3 = 2k 3/2 = k (3, 6) 3 = k(6) 3 = 6k 3/6 = k This table is NOT proportional because it has different constants.

OpenStudy (igreen):

Take two points from the third table: (3, 4.5) (5, 7.5) y = kx (3, 4.5) 4.5 = k(3) 4.5 = 3k 1.5 = k (5, 7.5) 7.5 = k(5) 7.5 = 5k 1.5 = k This table IS proportional because it has the same constant.

OpenStudy (igreen):

@ilovethemiddletvshow